OpenStudy (anonymous):

Debbie Anderson cuts a piece of wood measuring 5 1/8 inches into tow eaueal pieces. How long is each piece?

OpenStudy (lgbasallote):

simple. divide by 2 \[\huge 5 \frac 18 \div 2\] you'll get your answer do you need more help?

OpenStudy (nottim):

2 equal pieces?

OpenStudy (anonymous):

I come up with 2.5625. is this correct, and how do I convert it back to interger, proper fraction or mixed number?

OpenStudy (lgbasallote):

\[\Large 5 \frac 18 \div 2 \implies \frac{41}{8} \div 2 \implies \frac{41}{8} \times \frac 12 \implies \frac{41}{8 \times 2}\] do you follow?

OpenStudy (anonymous):

wouldn't that be 41/16?

OpenStudy (anonymous):

or is that the final answer?

OpenStudy (lgbasallote):

yup